OpenStudy (anonymous):

Paul, Raul, and Saul had a competition to see who could pick the greatest amount of strawberries in three minutes. Paul picked 1 5/8 pounds of strawberries, Raul 1 3/8 pounds, and Saul 1 11/16 pounds who came in second.

OpenStudy (anonymous):

First turn the mixed fractions into improper fractions and evaluate the result you get, arranging them from high to low...

OpenStudy (anonymous):

I'll do the one for paul\[1\frac{ 5 }{ 8 } \rightarrow \frac{ 13 }{ 8 }\]

OpenStudy (anonymous):

That's (1 x 8) + 5 = 13 Divide 13 by 8 (the denominator) See if you can attempt the others?

OpenStudy (anonymous):

1 3/8 = 11/8

OpenStudy (anonymous):

matttsheldon, that is wrong! Post your question in a separate thread. Thank you.

OpenStudy (anonymous):

Good work, dominick. That's for Raul. How about the one for Saul?

OpenStudy (anonymous):

7/112

OpenStudy (anonymous):

27/11

OpenStudy (anonymous):

Nice! So, we'll convert them to evaluate them each 13/8 = ?, 11/8 = ?, 27/11 = ?

OpenStudy (anonymous):

can you help me convert them

OpenStudy (anonymous):

Do you have a working calculator? Use it.

OpenStudy (anonymous):

how do you convert them

OpenStudy (anonymous):

Press 13 ÷ 8

OpenStudy (anonymous):

1.625

OpenStudy (anonymous):

Good, the others?

OpenStudy (anonymous):

1.375

OpenStudy (anonymous):

2.45

OpenStudy (anonymous):

So arrange them thus Paul = 1.625, Raul = 1.375 and Saul = 2.45 If we arrange them from lowest to highest, which of the boys come second?

OpenStudy (anonymous):

raul

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

My arrangement from here went this way 1.375, 1.625, 2.45 in that order

OpenStudy (anonymous):

Or if we should start from the highest picker, 2.45, 1.625, 1.375 Saul Paul Raul So, clearly Paul came in second.